How many electrons are necessary to produce a charge of -1.0 C? What is the mass of many electrons?
1
Expert's answer
2020-04-02T08:57:00-0400

Charge of one electron is -1.6*10-19 C, so to produce charge of -1.0 C the amount of

-1/(-1.6*10-19) = 6.25*1018 electrons is necessarry.

Mass of one electron is 9.1*10-31 kg, so the mass of the above amount of electrons is

6.25*1018*9.1*10-31 = 5.687*10-12 kg